# DISCO: Distribution-Aware Calibration for Object Detection with Noisy Bounding Boxes

### v2

@@ 3.3 Distribution-Aware Proposal Augmentation
Instead of using heuristic approaches such as selective search Uijlings et al. (2013) and edge box Zitnick and Dollár (2014), we propose to augment proposals with the modeled distribution, aiming to statistically cover more potential locations of objects. Firstly, we create a Guassian noise matrix $\mathbf{G}\in\mathbb{R}^{N^{\prime}\times 4}$ whose each element is sampled from $\mathcal{N}(0,1)$ to ensure randomness. Here $N^{\prime}$ is a hyperparameter that indicates the number of augmented proposals. Then, augmented proposals $\mathbf{P}^{i\prime}$ can be generated by
@@
where $\odot$ denotes element-wise multiplication and the operations here are all conducted in a broadcasting fashion. Following Equation 4 and 5, its classification scores $S^{i\prime}$ can be obtained. Subsequently, we incorporate the augmented proposals $\mathbf{P}^{i\prime}$ into $\mathbf{P}^{i}$, expressed as
$\mathbf{P}^{i*}\leftarrow\mathbf{P}^{i*}\oplus\mathbf{P}^{i\prime}\in\mathbb{R}^{(N^{i}+N^{\prime})\times 4},$ (10)
$S^{i}\leftarrow S^{i}\oplus S^{i\prime}\in\mathbb{R}^{N^{i}+N^{\prime}},$ (11)
where $\oplus$ indicates proposal-wise concatenation. To boost classification performance, the proposals with the highest classification score, which could contain representative category features, are collected to form a loss term for classification, formulated as
$s^{i}_{\star}=\max_{j}s^{i}_{j},\quad j=1,2,...,(N^{i}+N^{\prime}),$ (12)
@@
where $M$ is the number of proposal groups in a batch. Note that $\mathcal{L}_{\text{Aug}}$ does not be computed in the first-time DISCO. Finally, after proposal augmentation, we follow Equation 7 and 8 to model the spatial distribution of $\mathbf{P}^{i*}$ once again for a better representation of these proposals.
@@
RetinaNet 71.5 67.5 57.9 45.0 30.0 53.1 30.8 17.9 33.7 38.2 13.3 33.6 5.7 8.4 15.9 18.0
Co-teaching 75.4 70.6 60.9 43.7 30.5 54.9 30.5 17.3 34.0 39.1 11.5 31.4 4.2 6.4 13.1 16.4
SD-LocNet 75.7 71.5 60.8 43.9 30.0 54.5 30.3 17.5 33.6 38.7 11.3 30.3 4.3 6.0 12.7 16.6
FreeAnchor 73.0 67.5 56.2 41.6 28.6 53.1 28.5 16.6 32.2 37.0 10.4 28.9 3.3 5.8 12.1 14.9
KL Loss 75.8 72.7 64.6 48.6 31.0 54.3 32.4 18.0 34.9 39.5 12.1 36.7 3.7 6.2 13.0 17.4
OA-MIL 77.4 74.3 70.6 63.8 32.1 55.3 33.2 18.1 35.8 41.6 18.6 42.6 12.9 9.2 19.0 26.5
@@
Table 1: Benchmark results on VOC and COCO. Note that Clean-FasterRCNN is trained with clean annotations for reference. In noisy settings, the best results are marked in bold. Our DISCO can achieve state-of-the-art performance especially at high noise levels.
@@
Table 2: Experimental results on COCO at 10% and 30% noise levels. Our DISCO can still outperform OA-MIL and achieve state-of-the-art performance in additional noisy settings.
3.4 Distribution-Aware Box Refinement
where $\phi(\cdot)$ is a non-linear weighting function following Liu et al. (2022). To stabilize the early stage of training, the fusion strategy is conditional on the classification score $s^{i}_{\mu}$, and thus the $\phi(\cdot)$ is defined as
$\phi(s^{i}_{\mu})=\min((s^{i}_{\mu})^{\alpha},\beta),$ (15)
where $\alpha$ and $\beta$ are two hyperparameters. It means that the higher $s^{i}_{\mu}$ is, the more the model would reply on $\mu^{i}$. Finally, the refined $B^{i\prime}$ is used as supervision for the original proposals $\mathbf{P}^{i}$ to compute regression loss as
where $\operatorname{Dist}(\cdot,\cdot)$ is a predefined distance function for two bounding boxes Ren et al. (2015). It is worth noting that the first-time DISCO is ended with Equation 14 to obtain the refined ground-truth $B^{i\prime}$ for proposal re-assignment.
3.5 Distribution-Aware Confidence Estimation
To estimate the confidence of predicted bounding boxes, we integrate an estimator $\mathcal{E}(\cdot)$ into the DA head. Note that $\mathcal{E}(\cdot)$ comprises only one fully-connected layer. As excepted, the confidence $\mathbf{V}^{i}$ for the proposals $\mathbf{P}^{i}$ can be produced as
Different from He et al. (2019), we train the estimator with direct supervision of variance rather than implicit supervision of bounding boxes. Then, the estimated confidence (i.e., predicted variance) is used in Softer-NMS He et al. (2019) for a better inference-time process. Finally, the overall loss function is formed as
where $\gamma$ and $\lambda$ is two hyperparameters to down-weight $\mathcal{L}_{\text{Est}}$ and $\mathcal{L}_{\text{Aug}}$ respectively, and $\mathcal{L}_{\text{Cls}}$ is a cross-entropy classification loss for the original proposals $\mathbf{P}^{i}$ Ren et al. (2015).
@@ 3.5 Distribution-Aware Confidence Estimation
Table 3: Experimental results on the original VOC and COCO. Unlike OA-MIL, Our DISCO can still provide performance improvement even without manually introducing noise.
4 Experiments
4.1 Experimental Setup
Datasets.
Two large-scale image datasets are adopted in our experiments, including Pascal VOC 2007 Everingham et al. (2010) and MS-COCO 2017 Lin et al. (2014). Pascal VOC 2007 (VOC) is a standard dataset for object detection, consisting of $9,963$ images with $24,640$ box annotations. MS-COCO 2017 (COCO) is also a popular object detection benchmark, containing $328,000$ images of generic objects. Following Liu et al. (2022), noisy box annotations are simulated by perturbing clean ones at various noise levels which are set to $\{10\%,20\%,30\%,40\%\}$ for VOC and $\{20\%,40\%\}$ for COCO (see more details in Appendix A.1).
@@ 4.1 Experimental Setup
Hyper. Value $\text{AP}_{50}$bigstrut
$T$ 0.01 68.6
0.1 68.7
0.2 67.8
Hyper. Value $\text{AP}_{50}$bigstrut
$\alpha$ 3 68.3
5 68.7
7 68.2
Hyper. Value $\text{AP}_{50}$bigstrut
$\beta$ 0.7 68.1
0.8 68.7
0.9 67.3
Implementation Details.
Following Liu et al. (2022), we implement our method on FasterRCNN Ren et al. (2015) with ResNet-50 He et al. (2016) as the backbone. The idea of DISCO can be easily generalized to other frameworks and we choose to perform our experiments with FasterRCNN as it is widely adopted Kaur and Singh (2023). As a common practice, the model is trained with the “1×1$\times$1 ×” schedule Girshick et al. (2018). Hyperparameter selections of our method are detailed in Appendix A.2. Notably, all other training configurations are aligned with Liu et al. (2022) to ensure fairness.
Evaluation Metrics.
As commonly done, mean average precision (mAP@.5.5.5.5) and mAP@[.$[.5,95]$ are used for VOC and COCO respectively. Specifically, we report $\text{AP}_{50}$ for VOC and $\{\text{AP},\text{AP}_{50},\text{AP}_{75},\text{AP}_{S},\text{AP}_{M},\text{AP}_{L}\}$ for COCO.
4.2 Results and Discussions
We compare DISCO with the state-of-the-art methods of this task, including FasterRCNN Ren et al. (2015), Co-teaching Han et al. (2018), SD-LocNet Zhang et al. (2019a), KL Loss He et al. (2019), and OA-MIL Liu et al. (2022). Besides, the results of two one-stage methods are presented for a further comparison, including RetinaNet Lin et al. (2017) and FreeAnchor Zhang et al. (2019b). For reference, we also report the result of Clean-FasterRCNN, which is trained with clean annotations under the same setup.
Benchmark Results.
Benchmark results are reported in Table 1. It can be observed that noisy bounding box annotations significantly reduce the performance of vanilla object detectors like FasterRCNN, especially at high box noise levels. Moreover, Co-teaching and SD-LocNet can only marginally improve detection performance, showing that small-loss sample selection and sample weight assignment are not decent solutions for handling noisy bounding boxes. Besides, even with better label assignment, FreeAnchor still underperforms in such a challenging task. It is worth noting that KL Loss is a competitive method that also improves the interpretability of detectors. Moreover, OA-MIL adopts a MIL-based training strategy by iteratively constructing object-level bags, attaining better detection performance than the aforementioned methods. As shown in Table 1, our DISCO, which aims to calibrate the corrupt supervision signals caused by noisy bounding boxes, achieves state-of-the-art performance on these two benchmarks. Notably, it can significantly outperform the existing methods at high noise levels (i.e., $30\%$ and $40\%$), showing that our method is more robust to noisy bounding boxes. Specifically, compared with OA-MIL, DISCO attains $+1.5\text{AP}_{50}$ and $+4.9\text{AP}_{50}$ improvement on VOC at the $30\%$ and $40\%$ noise levels respectively. DISCO can also achieve $+2.6\text{AP}$, $+3.1\text{AP}_{50}$, and $+4.0\text{AP}_{75}$ improvement on COCO at the $40\%$ noise level.
Additional Evaluations.
To further demonstrate the effectiveness of our DISCO, We compare it with OA-MIL in more settings other than those included in Liu et al. (2022). The additional evaluations include two aspects: 1) Performance on COCO at $10\%$ and $30\%$ noise levels: Compared to OA-MIL, Our DISCO can still achieve state-of-the-art performance in additional noisy settings on COCO (see Figure 2), suggesting its flexibility for various noise levels; 2) Performance on the original VOC and COCO (i.e., the noise level is set to $0\%$): Without manually introducing noise, DISCO can still provide performance improvement on original datasets (see Figure 3), especially on VOC, showing that it has the potential to be generalized to real-world noisy scenarios.
Figure 5: Illustration of classification performance improvement. Left: Average classification scores of the positive proposals for corresponding categories. Right: Classification accuracy of the positive proposals. Compared to OA-MIL, DISCO can provide superior improvement for classification, which even approaches the results of training with clean annotations.
@@ 4.3 Ablation Studies
We conduct comprehensive ablation studies including component effectiveness and hyperparameter sensitivity to further verify DISCO’s performance. Due to space limitation, more ablation studies (e.g., backbone compatibility) are provided in Appendix B. Unless otherwise specified, the following experiments are all based on VOC at the $40\%$ noise level.
Component Effectiveness.
To investigate the effectiveness of three key techniques in DISCO (i.e., DA-Aug, DA-Ref, and DA-Est), we gradually integrate them into training. Note that the implementation of DA-Est is heavily based on DA-Ref thus it cannot be adopted independently. The experimental results are reported in Table 4, where we also list per-category performance for a detailed comparison. Notably, using only DA-Aug or DA-Ref can considerably contribute to performance improvement. DA-Ref seems to be more effective since it comes with refined ground-truths for better localization. Moreover, its detection performance can be further boosted when collaborating with DA-Aug or DA-Est. It is also worth noting that DA-Est can achieve $+4.1\text{AP}_{50}$ improvement by enhancing the robustness of detectors. Adopting all three techniques, our DISCO can attain superior detection performance in almost all categories, demonstrating the performance improvement of these components.
Hyperparameter Sensitivity.
Here we evaluate the sensitivity of $T$ of Equation 6 and $\alpha,\beta$ of Equation 15. The evaluations of other hyperparameters are provided in Appendix B.2. Note that we choose some moderate values rather than extreme ones to reasonably evaluate the sensitivity. As shown in Table 5, the temperature coefficient $T$ is relatively robust when set to 0.01 or 0.2. Tuning $T$ to a proper value can contribute to better performance. Moreover, the hyperparameters regulate the fusion of two bounding boxes (i.e., $\alpha$ and $\beta$) is also insensitive when varying within a moderate range, showing the effectiveness of our method.
@@ 4.4 Further Analysis
In this subsection, additional evidence and discussion are provided to further analyze the advantages of DISCO in classification, localization, and interpretability, respectively. Unless otherwise specified, the following experiments are all based on VOC at the $40\%$ noise level.
Classification Performance Improvement.
DA-Aug is used to generate proposals in the potential locations of objects for obtaining representative category features, by which classification performance can be boosted. The evidence is provided in Figure 5. It shows that noisy bounding box annotations can badly reduce the classification scores and the accuracy of foreground features. Notably, OA-MIL can also enhance classification performance. More importantly, DISCO provides superior improvement for classification, which even approaches the results of training with clean annotations.
Box Refinement for Better Localization.
To improve the localization capability of detectors, DA-Ref utilizes the modeled distributions of proposals for noisy box refinement (see Figure 6). Note that DISCO is performed twice in a training iteration and thus there are two successive refined boxes, where the first one is for proposal re-assignment and the second one acts as the supervision for regression. Since box refinement is very challenging when given only noisy ground-truths, it is natural that refined ones may be not exactly identical to real ones. Even so, the refined boxes can cover the objects more tightly than noisy ground-truths. Furthermore, the second-time refinement contributes to more precise ones, showing the effectiveness of our refinement strategy.
Interpretability with Confidence Estimation.
We introduce interpretability into box predictions in DISCO, aiming to enhance the robustness of detectors. The term “interpretability” is used to convey that our box predictions are interpretable, which is also adopted in He et al. (2019). This is implemented by estimating the confidence of each border of the predicted bounding boxes with the variance of the modeled distribution as its supervision. For an intuitive understanding, some qualitative results are represented in Figure 7. For a predicted border that deviates largely from the real one, DISCO could estimate a relatively large variance, indicating low confidence for this prediction. Such a crucial property enhances the practicability of DISCO in realistic scenarios.

### v3

@@ 3.2 Distribution-Aware Proposal Augmentation
Instead of using heuristic approaches such as selective search [Uijlings et al.(2013)Uijlings, Van De Sande, Gevers, and Smeulders] and edge box [Zitnick and Dollár(2014)], we propose to augment proposals with the modeled distribution, aiming to statistically cover more potential locations of objects. Firstly, we create a Guassian noise matrix $\mathbf{G}\in\mathbb{R}^{N^{\prime}\times 4}$ whose each element is sampled from $\mathcal{N}(0,1)$ to ensure randomness. Here $N^{\prime}$ is a hyperparameter that indicates the number of augmented proposals. Then, augmented proposals $\mathbf{P}^{i\prime}$ can be generated by
@@
where $\odot$ denotes element-wise multiplication and the operations here are all conducted in a broadcasting fashion. Following Equation 3, its classification scores $S^{i\prime}$ can be obtained. Subsequently, we incorporate the augmented proposals $\mathbf{P}^{i\prime}$ into $\mathbf{P}^{i}$, expressed as
$\mathbf{P}^{i*}\leftarrow\mathbf{P}^{i*}\oplus\mathbf{P}^{i\prime}\in\mathbb{R}^{(N^{i}+N^{\prime})\times 4},\quad S^{i}\leftarrow S^{i}\oplus S^{i\prime}\in\mathbb{R}^{N^{i}+N^{\prime}},$ (7)
where $\oplus$ indicates proposal-wise concatenation. To boost classification performance, the proposals with the highest classification score $s^{i}_{\star}=\max_{j}s^{i}_{j}$, which could contain representative category features, are collected to form a loss term for classification, formulated as
@@
where $M$ is the number of proposal groups in a batch. Note that $\mathcal{L}_{\text{Aug}}$ does not be computed in the first-time DISCO. Finally, after proposal augmentation, we follow Equation 5 to model the spatial distribution of $\mathbf{P}^{i*}$ once again for a better representation of these proposals.
3.3 Distribution-Aware Box Refinement
where $\phi(\cdot)$ is a non-linear weighting function. To stabilize the early stage of training, the fusion strategy is conditional on the classification score $s^{i}_{\mu}$, and thus the $\phi(\cdot)$ is defined as
$\phi(s^{i}_{\mu})=\min((s^{i}_{\mu})^{\alpha},\beta),$ (10)
where $\alpha$ and $\beta$ are two hyperparameters. It means that the higher $s^{i}_{\mu}$ is, the more the model would rely on $\mu^{i}$. Finally, the refined $B^{i\prime}$ is used as supervision for the original proposals $\mathbf{P}^{i}$ to compute regression loss as
where $\operatorname{Dist}(\cdot,\cdot)$ is a predefined distance function for two bounding boxes [Ren et al.(2015)Ren, He, Girshick, and Sun]. It is worth noting that the first-time DISCO is ended with Equation 9 to obtain the refined ground-truth $B^{i\prime}$ for proposal re-assignment.
3.4 Distribution-Aware Confidence Estimation
To estimate the confidence of predicted bounding boxes, we integrate an estimator $\mathcal{E}(\cdot)$ into the DA head. Note that $\mathcal{E}(\cdot)$ comprises only one fully connected layer. As excepted, the confidence $\mathbf{V}^{i}$ for the proposals $\mathbf{P}^{i}$ can be produced as
Different from [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang], we train the estimator with direct supervision of variance rather than implicit supervision of bounding boxes. Then, the estimated confidence (i.e., predicted variance) is used in Softer-NMS [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang] for a better inference-time process. Finally, the overall loss function is formed as
where $\gamma$ and $\lambda$ is two hyperparameters to down-weight $\mathcal{L}_{\text{Est}}$ and $\mathcal{L}_{\text{Aug}}$ respectively, and $\mathcal{L}_{\text{Cls}}$ is a cross-entropy classification loss for the original proposals $\mathbf{P}^{i}$ [Ren et al.(2015)Ren, He, Girshick, and Sun].
@@ 3.4 Distribution-Aware Confidence Estimation
RetinaNet [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár] 71.5 67.5 57.9 45.0 30.0 53.1 30.8 17.9 33.7 38.2 13.3 33.6 5.7 8.4 15.9 18.0
Co-teaching [Han et al.(2018)Han, Yao, Yu, Niu, Xu, Hu, Tsang, and Sugiyama] 75.4 70.6 60.9 43.7 30.5 54.9 30.5 17.3 34.0 39.1 11.5 31.4 4.2 6.4 13.1 16.4
SD-LocNet [Zhang et al.(2019a)Zhang, Yang, and Feng] 75.7 71.5 60.8 43.9 30.0 54.5 30.3 17.5 33.6 38.7 11.3 30.3 4.3 6.0 12.7 16.6
FreeAnchor [Zhang et al.(2019b)Zhang, Wan, Liu, Ji, and Ye] 73.0 67.5 56.2 41.6 28.6 53.1 28.5 16.6 32.2 37.0 10.4 28.9 3.3 5.8 12.1 14.9
KL Loss [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang] 75.8 72.7 64.6 48.6 31.0 54.3 32.4 18.0 34.9 39.5 12.1 36.7 3.7 6.2 13.0 17.4
OA-MIL [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang] 77.4 74.3 70.6 63.8 32.1 55.3 33.2 18.1 35.8 41.6 18.6 42.6 12.9 9.2 19.0 26.5
@@
Table 1: Benchmark results on VOC and COCO. Note that Clean-FasterRCNN is trained with clean annotations for reference. In noisy settings, the best results are marked in bold.
4 Experiments
To ensure fairness, our experimental setup is all aligned with [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang]. Specifically, two large-scale detection datasets, i.e., Pascal VOC 2007 (VOC) [Everingham et al.(2010)Everingham, Van Gool, Williams, Winn, and Zisserman] and MS-COCO 2017 (COCO) [Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick], are adopted in the experiments. Noisy box annotations are simulated by perturbing clean ones at various noise levels which are set to $\{10\%,20\%,30\%,40\%\}$ for VOC and $\{20\%,40\%\}$ for COCO. We report $\text{AP}_{50}$ for VOC and $\{\text{AP},\text{AP}_{50},\text{AP}_{75},\text{AP}_{S},\text{AP}_{M},\text{AP}_{L}\}$ for COCO. Due to the space limit, more details are included in Appendix A.
@@ 4 Experiments
Table 2: Left: Experimental results on COCO at 10% and 30% noise levels. DISCO can still achieve SOTA performances. Right: Experimental results on the original VOC and COCO. DISCO can provide performance improvement without manually introducing noise.
4.1 Results and Discussions
We compare DISCO with the SOTA methods of this task, including FasterRCNN [Ren et al.(2015)Ren, He, Girshick, and Sun], Co-teaching [Han et al.(2018)Han, Yao, Yu, Niu, Xu, Hu, Tsang, and Sugiyama], SD-LocNet [Zhang et al.(2019a)Zhang, Yang, and Feng], KL Loss [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang], and OA-MIL [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang]. Besides, the results of two one-stage methods are presented for a further comparison, including RetinaNet [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár] and FreeAnchor [Zhang et al.(2019b)Zhang, Wan, Liu, Ji, and Ye]. For reference, we also report the result of Clean-FasterRCNN, which is trained with clean annotations under the same setup.
Benchmark Results. Benchmark results are reported in Table 1. It can be observed that noisy bounding box annotations significantly reduce the performance of vanilla object detectors like FasterRCNN, especially at high box noise levels. Moreover, Co-teaching and SD-LocNet can only marginally improve detection performance, showing that small-loss sample selection and sample weight assignment are not decent solutions for handling noisy bounding boxes. Besides, even with better label assignment, FreeAnchor still underperforms in such a challenging task. It is worth noting that KL Loss is a competitive method that also improves the interpretability of detectors. Moreover, OA-MIL adopts a MIL-based training strategy by iteratively constructing object-level bags, attaining better detection performance than the aforementioned methods. Our DISCO, which aims to calibrate the corrupt supervision signals caused by noisy bounding boxes, achieves SOTA performance on these two benchmarks. Notably, it can significantly outperform the existing methods at high noise levels (i.e., $30\%$ and $40\%$), showing that our method is more robust to noisy bounding boxes. Specifically, compared with OA-MIL, DISCO attains $+1.5\text{AP}_{50}$ and $+4.9\text{AP}_{50}$ improvement on VOC at the $30\%$ and $40\%$ noise levels respectively. DISCO can also achieve $+2.6\text{AP}$, $+3.1\text{AP}_{50}$, and $+4.0\text{AP}_{75}$ improvement on COCO at the $40\%$ noise level.
@@ 4.1 Results and Discussions
Additional Evaluations. To further demonstrate the effectiveness of DISCO, We compare it with OA-MIL in more settings other than those included in [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang]. The additional evaluations include two aspects (see Table 2): 1) Performance on COCO at $10\%$ and $30\%$ noise levels: Compared to OA-MIL, DISCO can still achieve SOTA performance in additional noisy settings on COCO, suggesting its flexibility for various noise levels; 2) Performance on the original VOC and COCO (i.e., the noise level is set to $0\%$): Without manually introducing noise, DISCO can still provide performance improvement on original datasets, especially on VOC, showing that it has the potential to be generalized to real-world noisy scenarios.
Figure 4: Illustration of classification performance improvement. Left: Average classification scores of the positive proposals for corresponding categories. Right: Classification accuracy of the positive proposals. Compared to OA-MIL, DISCO provides better improvement for classification, which even approaches the results of training with clean annotations.
@@ 4.2 Ablation Studies
To investigate the effectiveness of three key techniques in DISCO, we gradually integrate them into training. Note that the implementation of DA-Est is heavily based on DA-Ref thus it cannot be used independently. The experimental results are reported in Table 3, which are based on VOC at the $40\%$ noise level. Notably, using only DA-Aug or DA-Ref can considerably contribute to performance improvement. DA-Ref seems to be more effective since it comes with refined ground-truths for better localization. Moreover, its detection performance can be further boosted when collaborating with DA-Aug or DA-Est. It is also worth noting that DA-Est can achieve $+4.1\text{AP}_{50}$ improvement by enhancing the robustness of detectors. Using all three techniques, our DISCO can attain superior detection performance in almost all categories. Because of the space limit, more ablation studies such as hyperparameter sensitivity and backbone compatibility are provided in Appendix B.
Figure 5: (a) Qualitative results of box refinement in DISCO. Real ground-truths and noisy ground-truths are marked in orange and blue. Refined bounding boxes produced by the first-/second-time DISCO are indicated in dotted/solid red. The first-time refined boxes can cover the objects more tightly than noisy ground-truths, and the second-time refinement can further contribute to more precise ones. (b) Qualitative results of interpretability in DISCO. We randomly choose an assigned proposal (yellow) per image to report its estimated variances. Real ground-truths and noisy ground-truths are marked in orange and blue. Note that the variance is scaled by the width and height for clarity. With the proposed DA-Est, DISCO can estimate reasonable variances for each border of box prediction.
@@ 4.3 Further Analysis
In this subsection, additional evidence and discussion are provided to further analyze the advantages of DISCO in classification, localization, and interoperability. Unless otherwise specified, the following experiments are all based on VOC at the $40\%$ noise level.
Classification Performance Improvement. In DISCO, DA-Aug is used to generate proposals in the potential locations of objects for obtaining representative category features, by which classification performance can be boosted. The evidence for advantage is provided in Figure 4. It shows that noisy bounding box annotations can badly reduce the classification scores and the accuracy of foreground features. Notably, OA-MIL [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang] can also enhance classification performance. More importantly, DISCO provides superior improvement for classification, which even approaches the results of training with clean annotations.
Box Refinement for Better Localization. To improve the localization capability of detectors, DA-Ref utilizes the modeled distributions of proposals for noisy box refinement (see Figure 5(a)). Note that DISCO is performed twice in a training iteration and thus there are two successive refined boxes, where the first one is for proposal re-assignment and the second one acts as the supervision for regression. Since box refinement is very challenging when given only noisy ground-truths, it is natural that refined ones may be not exactly identical to real ones. Even so, the refined boxes can cover the objects more tightly than noisy ground-truths. Furthermore, the second-time refinement contributes to more precise ones, showing the effectiveness of our refinement strategy.
Interpretability with Confidence Estimation. We introduce interpretability into box predictions in DISCO, aiming to enhance the robustness of detectors. The term "interpretability" is used to convey that our box predictions are interpretable, which is also adopted in [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang]. This is implemented by estimating the confidence of each border of the predicted bounding boxes with the variance of the modeled distribution as its supervision (see Figure 5(b)). For a predicted border that deviates largely from the real one, DISCO could estimate a relatively large variance, indicating low confidence for this prediction. Such a crucial property enhances the practicability of DISCO in realistic scenarios.